OpenStudy (anonymous):

An item is priced at $14.99. If the sales tax is 6%, what does the item cost including sales tax? $17.39 $20.99 $15.89 $15.05 @coniquee

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Thanks... here's another one.. The value of Peter's house has risen 10% a year for the last 3 years. If the value of Peter's home 3 years ago was $90,000, what is the value now? $108,900.0000000000 $117,000 $99,000.0000000000 $119,790.0000000000

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

Hold on sorry

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

it says it's D

OpenStudy (anonymous):

Ugh auto correct its D

OpenStudy (anonymous):

Well you got it

OpenStudy (anonymous):

A jacket usually sells for $69.00. If the jacket is 40% off, and sales tax is 6%, what is the total price of the jacket, including tax? $44.13 $41.40 $73.27 $43.88 Last 2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Could you explain how you got that please? :)

OpenStudy (anonymous):

Okay 69.00 -40% = 27.60 + 6%= 43.88

OpenStudy (anonymous):

Ohh I see. thank you!! I have 3 more if thats ok

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

A toaster oven usually sells for $61. If the toaster oven is 20% off, and sales tax is 5%, what is the total price of the toaster oven, including tax? $51.24 $64.05 $48.80 $61.74

OpenStudy (anonymous):

51.24

OpenStudy (anonymous):

Could you show me how you got that ? :)

OpenStudy (anonymous):

wait i think i know $48.80 + ($48.80 × 0.05) = $51.24 right?